if cost rate is rs.2 per kwh then 500 W kettle operate for 30 min will cost 
ans . 0.50 Rs

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
Totall units = 500 *1/2 = 250 watt hour = o.25 KWH
 
Hence total cost = 0.25 * 2 = 0.50 Rs